### v2

# SAE-V: Interpreting Multimodal Models for Enhanced Alignment

Abstract
With the integration of image modality, the semantic space of multimodal large language models (MLLMs) is more complex than text-only models, making their interpretability more challenging and their alignment less stable, particularly susceptible to low-quality data, which can lead to inconsistencies between modalities, hallucinations, and biased outputs. As a result, developing interpretability methods for MLLMs is crucial for improving alignment quality and efficiency. In text-only LLMs, Sparse Autoencoders (SAEs) have gained attention for their ability to interpret latent representations. However, extending SAEs to multimodal settings presents new challenges due to modality fusion and the difficulty of isolating cross-modal representations. To address these challenges, we introduce SAE-V, a mechanistic interpretability framework that extends the SAE paradigm to MLLMs. By identifying and analyzing interpretable features along with their corresponding data, SAE-V enables fine-grained interpretation of both model behavior and data quality, facilitating a deeper understanding of cross-modal interactions and alignment dynamics. Moreover, by utilizing cross-modal feature weighting, SAE-V provides an intrinsic data filtering mechanism to enhance model alignment without requiring additional models. Specifically, when applied to the alignment process of MLLMs, SAE-V-based data filtering methods could achieve more than 110% performance with less than 50% data. Our results highlight SAE-V’s ability to enhance interpretability and alignment in MLLMs, providing insights into their internal mechanisms. ***Our codebase and model are released at Github and Huggingface.
Machine Learning, ICML
Figure 1: Operational Dynamics of SAE-V Based Data Filtering Method. SAE-V encodes and interprets the representation inside MLLM during alignment and inference time. Based on this representation, we could reveal the modality gap within the data, and improve the alignment process through the selection of modality-fused, high-quality data. This pipeline performs data filtering without requiring additional models, relying instead on MLLM itself to prioritize high-value data effectively.
1 Introduction
With the development and success of large language models (LLMs) (Dubey et al., 2024; Achiam et al., 2023), researchers have begun to introduce visual understanding to these models, thereby extending their operational scope from language to a mix of vision and language, resulting in the creation of powerful multimodal large language models (MLLMs) (Alayrac et al., 2022; Liu et al., 2024a; Team et al., 2024; Team, 2024). To enhance the multimodal understanding capabilities of MLLMs, the research community has explored various architectures, including using individual image/text encoders to encode cross-modal information into a joint representation space (Zhang et al., 2023; Liu et al., 2024a; Zhu et al., 2024; Wu et al., 2024c) and leveraging image tokenizers to transform all inputs into a unified token sequence (Team, 2024; Xie et al., 2024; Wu et al., 2024a; Wang et al., 2024). Despite the difference in the architectures of these models, their essential goal is the same: Fuse the text and image representation space into a joint multimodal semantic space.
As MLLMs continue to scale up in both size and capability, their interpretability and controllability remain a significant challenge (Zhang & Zhu, 2018; Ben Melech Stan et al., 2024). Currently, mechanistic interpretability techniques such as circuit analysis (Olsson et al., 2022) and dictionary learning with sparse autoencoders (Huben et al., 2024) are the most widely recognized approaches to interpreting LLMs. However, their application to MLLMs, especially in the context of cross-modal integration, has been limited. There is a pressing need for specialized tools and frameworks that can unravel the intricate workings of MLLMs.
Moreover, current interpretability efforts are focused mainly on interpreting models, rather than applying this interpretability to real alignment situations, which also makes it difficult to evaluate these methods effectively. Top-down approaches, such as Representation Engineering (Zou et al., 2023) and activation steering (Turner et al., 2023; Rimsky et al., 2024), can directly evaluate the control effects of interpretability methods through control or unlearning techniques. However, for bottom-up methods like circuit analysis (Wang et al., 2023), sparse autoencoders (SAEs) (Huben et al., 2024), and cross-coders, effective evaluation methods beyond loss reduction are limited. Based on the previous discussion, can we propose a bottom-up multimodal interpretability approach that can directly enhance the alignment process?
In this work, we developed SAE-V, a mechanistic interpretability framework for MLLMs that extends the SAE paradigm to MLLMs. These tools are then applied to interpret the training process of transitioning from LLMs to MLLMs, as well as the process of enhancing the multimodal capabilities of MLLMs. Furthermore, utilizing the interpretable features of SAE-V models and their relationship to MLLM capabilities, we designed a data filtering metric based on SAE-V. This metric can filter out data that hinder the development of multimodal understanding, achieving stronger alignment with a smaller dataset. Overall, our work makes the following contributions:
Figure 2: The interpretability and data filtering pipeline of SAE-V. SAE-V is trained to encode MLLM activations into sparse, interpretable features. We first acquire the cross-modal weight of these features via SAE-V models, then reversely score the given data by the weighted average of each feature’s score. In this way, we provide an intrinsic data filtering tool by eliciting MLLM’s latent representation of these data.
  • Multimodal interpretability tool We developed mechanistic interpretability tools for MLLMs based on previous attempts on LLMs and trained corresponding SAE-V models. We demonstrated that SAE-V models trained on MLLMs can effectively extract interpretable features, and SAE-V models can be transferred to the corresponding LLMs. Specifically, the reconstruction loss of our SAE-V models trained on MLLMs is 38.3% and 50.6% compared to the SAE model when applied to MLLMs and LLMs, respectively.
  • Interpreting Multimodal Alignment Process We utilized SAE-V to study the feature distribution throughout the alignment process. We discovered that the feature distribution of SAE-V corresponds to the MLLM’s performance on multimodal understanding tasks.
  • Filtering metric to improve alignment Based on the previous investigation with SAE-V, we developed a metric to filter multimodal datasets and acquire high-quality data, therefore improving alignment quality and efficiency. Experiments demonstrate that our filtering tool achieves more than 110% performance compared to the full dataset while using 50% less data, underscoring the efficiency and effectiveness of SAE-V.
2 Methodology
In this section, we present our method to train, evaluate, and apply SAE-V to interpret MLLMs and multimodal data.
2.1 Preliminary: Sparse Autoencoder Paradigm
We adopt SAE-V (denoted as $\mathcal{S}_{\theta}$) architecture from the methodology proposed in (Bricken et al., 2023), which comprises an encoder and a feature dictionary $\mathcal{F}_{\theta}\text{:}\ \{\bm{f}_{k}\}_{k=1}^{\bm{n}}$ as a decoder. Let the input be denoted as $H\in\mathbb{R}^{\bm{l}\times\bm{m}}$, where $\bm{l}$ donates the number of input tokens and $\bm{m}$ donates the shape of hidden state token, the hidden state of a specific layer of a MLLM $\mathcal{M}_{\theta}$. The SAE-V encoding operation $\mathcal{S}_{\theta}(\cdot)$ is defined as
$Z=\text{ReLU}(H\times W_{\text{enc}}+b_{enc}),$ (1)
where $Z\in\mathbb{R}^{\bm{l}\times\bm{n}}$ is the feature activation of the input. The reconstruction loss of $\mathcal{S}_{\theta}$ donates as
$\mathcal{L}_{R}={||H-Z\times(\bm{f}_{1},\bm{f}_{2},\dots,\bm{f}_{n})^{\top}||}_{2}^{2}.$ (2)
The training loss is defined by
$\mathcal{L}=\mathcal{L}_{R}+\lambda\mathcal{L}_{1},$ (3)
where $\mathcal{L}_{1}=||Z||_{1}$ adds a sparsity constraint to the learned features and $\lambda$ is a hyperparameter controlling the level of sparsity. The training results could also quantized by incorporating an additional sparsity constraint via $\mathcal{L}_{0}=||Z||_{0}$, which counts the number of nonzero elements in the learned features $Z$.
Algorithm 1 Cosine similarity score Ranking
  Require: Text token vocabulary: $\mathcal{T}$; vision token vocabulary: $\mathcal{V}$; multimodal dataset $\mathcal{D}=\{\bm{d}_{i}\}_{i=1}^{p}$; MLLM parameterized by $\mathcal{M_{\theta}}$; SAE-V model $\mathcal{S_{\theta}}$; features of SAE-V model $\mathcal{F_{\theta}}\text{:}\ \{\bm{f}_{k}\}_{k=1}^{n}$; activation bound: $\delta$; cosine similarity function C (Equation 7)
  Ensure: Ranked data $\mathcal{D}_{R}$
  Stage 1: Collect Feature Activation Token
  Initialize activated token set of features $\mathcal{A}_{k}\leftarrow\varnothing$
  $\mathcal{D}_{s}\leftarrow\textit{Sample}(\mathcal{D})$
  for each $\bm{d}_{i}\in\mathcal{D}_{s}$ do
     $H_{i}\leftarrow\mathcal{M_{\theta}}(\bm{d}_{i})$
     $Z_{i}\leftarrow\mathcal{S_{\theta}}(H_{i})$
     for each $\bm{f}_{k}\in\mathcal{F_{\theta}}$ do
        $\mathcal{A}_{k}\leftarrow\mathcal{A}_{k}\cup\{\bm{h}_{j}\text{:}\ \bm{h}_{j}\in H_{i},\bm{z}_{j}=\bm{e}_{j}Z_{i},\bm{z}_{jk}>\delta\}$
     end for
  end for
  Stage 2: Compute Cross-modal Weight
  Initialize cross-modal weight of features $\mathcal{\omega}_{k}\leftarrow 0$
  for each $\bm{f}_{k}\in\mathcal{F_{\theta}}$ do
     $\mathcal{\omega}_{k}\leftarrow\textit{C}(\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{T}),\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{V}))$
  end for
  Stage 3: Rank Dataset by Cross-modal Weight
  Initialize cross-modal score of data $\bm{s}_{i}\leftarrow 0$
  for each $\bm{d}_{i}\in\mathcal{D}$ do
     $Z_{i}\leftarrow\mathcal{S_{\theta}}(\mathcal{M_{\theta}}(\bm{d}_{i}))$
     $F_{i}\leftarrow\{\bm{f}_{k}\text{: }\exists\ z_{j}\in Z_{i},\ \mathrm{s.t.}\ z_{j}\ \text{activates }\bm{f}_{k}\}$
     $\bm{s}_{i}\leftarrow\sum_{\bm{f}_{k}\in F_{i}}\omega_{k}$
  end for
  $\mathcal{D}_{R}\leftarrow\textit{Sort}(\mathcal{D},\{\bm{s}_{i}\}_{i=1}^{n})$
2.2 Interpreting Multimodal Data with SAE-V
It has been previously demonstrated (Gao et al., 2025; Huben et al., 2024) that SAE can be employed to interpret how LLMs encode semantic information from these models. This feature motivates us to apply SAE-V to assess the quality of the data and thus facilitate data filter for alignment.
We adopt a cosine similarity score ranking algorithm for data filtering (shown in Algorithm 1). Let the multimodal training dataset be donated as $\mathcal{D}=\{\bm{d}_{i}\}_{i=1}^{p}$, where $\bm{d}_{i}=(u_{\bm{1}},u_{\bm{2}},\dots,u_{\bm{l}})$, $u_{\bm{j}}\in\mathcal{T}\lor u_{\bm{j}}\in\mathcal{V},\bm{j}=\bm{1},\bm{2},\dots,\bm{l}$, is a sequence of tokens from text vocabulary $\mathcal{T}$ and tokens from vision vocabulary $\mathcal{V}$. We acquire feature activation token $\bm{z}_{j}$ by MLLM forward and Equation 1, i.e.
$H_{i}$ $=\mathcal{M}_{\theta}(\bm{d_{i}})$ (4)
$Z_{i}$ $=\mathcal{S}_{\theta}(H_{i}),$ (5)
$\bm{z}_{j}$ $=\bm{e}_{j}Z_{i},$ (6)
$\text{where}\ \bm{e}_{j}$ $=(0,0,\ldots,\underbrace{1}_{j\text{-th position}},\ldots,0).$
We define a SAE-V feature $\bm{f}_{k}$ is activated on $\bm{z}_{j}\ \textit{if}\ \bm{z}_{jk}>\delta$, where $\delta$ is activation bound. Correspondingly, we state that $\bm{f}_{k}$ is activated on $\bm{d}_{i}\ \textit{if}\ \exists\ \bm{z}_{j}\in Z_{i},\ \mathrm{s.t.}\ z_{j}\ \text{activates}\ \bm{f}_{k}$.
Our algorithm 1 consists of three stages: (1) Collecting feature activation tokens from dataset, (2) Computing cross-modal weight of SAE-V features, and (3) Ranking dataset by cross-modal weight.
  1. 1.
    Feature Activation Token Collecting We first sample a small subset $\mathcal{D}_{S}$ of the training dataset $\mathcal{D}$ and input these samples into the MLLM to obtain hidden states $H$. These hidden states are then fed into the SAE-V encoder to extract feature activations defined as Equation 4. For each feature, we collect its hidden state tokens thereby obtaining a sample of feature activation tokens across the dataset.
  2. 2.
    SAE-V Feature Weighting For each feature $\bm{f}_{k}$, we identify its top-K hidden state text tokens $\bm{t}=\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{T})$ and top-K hidden state vision tokens $\bm{v}=\textit{TopK}(\mathcal{A}_{k}\cap\mathcal{V})$, where the top-K is ranked by the activation value $\bm{z}_{jk}$ of the token. We then compute the cosine similarity between the two lists of tokens, donating the cross-modal weight of feature $\bm{f}_{k}$ as
    $Cosine(\bm{t},\bm{v})=\mathbb{E}_{i,j}\frac{\bm{t}_{i}\cdot\bm{v}_{j}}{||\bm{t}_{i}||||\bm{v}_{j}||},$ (7)
    which represents the capability of the feature to capture multimodal information within data.
  3. 3.
    Data Ranking Using the weighted features of SAE-V model, we score the entire training dataset. The cosine similarity score of each piece of data is defined as the sum of the cosine similarity scores of its activating features. We rank the data set by the score and the resulting cosine similarity score order allows us to filter data that are better aligned with the structures of multimodal semantic information.
We present our experiments and results in Section 4, demonstrating that our cosine similarity score ranking method can effectively filter high-quality data from the training data set.
3 Interpretability Analysis with SAE-V
Figure 3: Examples of interpretable features discovered by SAE-V. We presented examples of interpretable SAE-V features on LLaVA-NeXT-7B that demonstrate cross-modal semantic consistency. (a) Feature #44031 exhibits strong activation for ‘Doberman dogs’ across both text and image modalities, showing SAE-V’s ability to identify specific concepts with concrete physical meanings. (b) Feature #11105 captures the abstract concept of ‘Symmetry’ across different modalities, activating for various symmetry patterns including left-right, top-bottom, and central symmetry, with activation regions precisely aligned with symmetrical elements in images. These examples illustrate that SAE-V can discover features representing both concrete entities and abstract concepts that maintain consistent semantic meaning across modalities, going beyond what traditional probing methods on raw activations can achieve.
In this section, we conduct experiments on the SAE-V paradigm, aiming to demonstrate the capability and transferability of SAE-V model. We also performed experiment to prove the effectiveness of our SAE-V-based data interpreting tool from the inference side.
3.1 Training and Evaluating SAE-V Model
We trained a series of SAE and SAE-V models on MLLMs and their base LLMs. We evaluated the performance of these models, and the results demonstrated that SAE-V model is capable of interpreting MLLMs and that SAE-V model trained on MLLM can be effectively transferred to its original LLM.
3.1.1 Experiment Setup
Datasets
For text-only and multimodal situations, we selected the Pile (Gao et al., 2020) and Obelics (Laurençon et al., 2023) datasets separately. Specifically, we sampled 100K data from each dataset as the train set and 10K data as the test set. The Pile is a diverse language modeling dataset for LLM pretraining, and Obelics is a massive interleaved image-text dataset for MLLM pretraining. These two datasets are widely recognized in various pretraining and interpretability works (Black et al., 2022; Biderman et al., 2023; Huben et al., 2024; Team, 2024).
Models
We selected two generic MLLMs, LLaVA-NeXT models (including LLaVA-NeXT-Mistral-7B, LLaVA-NeXT-Vicuna-7B and LLaVA-NeXT-Vicuna-13B) (Liu et al., 2024a) and Chameleon-7B (Team, 2024), as our target models. These models represent two distinct architectures, and testing our method on them can demonstrate that our method is applicable to different architectures.
Additionally, we also studied Anole-7B (Chern et al., 2024) and Mistral-7B (Jiang et al., 2023) to compare the behavior of SAE and SAE-V models before and after fine-tuning, specifically the transitioning fine-tuning from LLM to MLLM. Anole-7B is a variant of Chameleon-7B, with its image generation capability unlocked, while Mistral-7B is the base LLM of LLaVA-NeXT-7B.  [^2] We present the detailed training setup and hyper-parameters in Appendix A.1.
Evaluation Metrics
To evaluate the performance of SAE-V models, we use two key metrics: $\mathcal{L}_{0}=||\bm{z}||_{0}$ where $\bm{z}$ is defined in Equation 6 and reconstruction loss $\mathcal{L}_{R}$ in Equation 2. $\mathcal{L}_{0}$ quantifies the number of activated features, reflecting the method’s ability to extract interpretable features, while reconstruction loss measures the method’s activation reconstruction capability compared with the model output, indicating the method’s accuracy in giving interpretations.
3.1.2 Experiment Result
Model Method $\mathcal{L}_{0}$
Chameleon-7B SAE 24757.6
SAE-V 50.1
Anole-7B SAE 62.1
SAE-V 50.1
LLaVA-NeXT-Mistral-7B SAE 94.5
SAE-V 192.5
LLaVA-NeXT-Vicuna-7B SAE 3162.96
SAE-V 585.64
LLaVA-NeXT-Vicuna-13B SAE 128.56
SAE-V 193.63
Table 1: The $\mathcal{L}_{0}$ metric of SAE and SAE-V models. $\mathcal{L}_{0}$ indicates the sparsity cost (average activated feature number). The results vary significantly across models due to their architectural differences. For Anole-7B and Chameleon-7B, SAE-V models maintain lower $\mathcal{L}_{0}$, suggesting more efficient feature utilization. However, LLaVA-NeXT models shows a contrary pattern with SAE-V requiring higher feature activation than SAE. We propose that extra activated features of SAE-V model are introduced by extra vision tokens in multimodal data. Notably, Chameleon-7B and LLaVA-NeXT-Vicuna-7B with SAE model exhibits an unusually high sparsity cost, attributed to multiple unseen vision tokens in the inference stage.
Figure 4: Reconstruction capability of SAE and SAE-V models. Each section compares the metrics of zero (set all activations as zero), SAE model, SAE-V model, and the Original reference state. SAE-V model consistently demonstrates superior reconstruction performance across all tested models. In Chameleon-7B and Anole-7B, SAE performs worse than the zero baseline, which indicates that SAE trained in text data fails to capture interpretable features in these MLLMs.
Capability of SAE-V Model
We compare the performance of SAE-V and SAE on different multimodal models. The $\mathcal{L}_{0}$ (shown in Table 1) varies significantly across the three models. For LLaVA-NeXT-7B, the $\mathcal{L}_{0}$ of SAE-V is much higher than that of SAE. For Chameleon-7B, SAE-V performs normally, whereas the $\mathcal{L}_{0}$ of SAE is abnormally high, indicating that SAE fails to extract sparse features. We suppose that the failure is attributed to a large number of unseen vision tokens for SAE during the inference stage. For Anole-7B, the $\mathcal{L}_{0}$ of SAE and SAE-V are nearly identical. The reconstruction loss (shown in Figure 4) of SAE-V is lower than SAE and is closer to the original activation, demonstrating that SAE-V behaves much better at reconstructing original activation than SAE across all three models. The results indicate that SAE-V outperforms SAE in terms of capability.
Figure 5: Reconstruction performance of SAE and SAE-V. The x-axis shows different models and task configurations, text indicates text-only task, and text & vision indicates multimodal task. The colored bars represent five experiment groups (zero activation, SAE of Mistral-7B, SAE of LLaVA-NeXT-7B trained with the Pile, SAE of LLaVA-NeXT-7B trained with Obliecs, original performance). Across various settings, SAE-V consistently demonstrates superior transferability compared to SAE and achieves reconstruction loss close to the original performance, the maximum relative gap being 67.28%. SAE based on Mistral-7B and SAE based on LLaVA-NeXT-7B achieves nearly the same loss in all tasks and models, indicating the equivalence of training SAE with MLLM and its base LLM.
Transferability of SAE-V Model
The transferability of SAEs between foundation models and instruction-tuned models has been extensively investigated in text-only contexts (Kissane et al., 2024; Kutsyk et al., 2024; Gallifant et al., 2025), as it demonstrates whether SAEs can capture universal semantic features within LLMs. Similarly, the transferability from MLLMs to corresponding LLMs serves as a critical metric for the quality of features learned by SAE-V. We compared the reconstruction performance of SAE-V model trained on LLaVA-NeXT-7B and SAE model to prove that SAE-V model trained on MLLMs can generalize to its base LLM. The findings (shown in Figure 5) indicate that across different settings, SAE-V model consistently achieves the best performance. Moreover, when trained on both MLLM and LLM, SAE model exhibits nearly identical reconstruction loss values, showing its robust transferability.
These results highlight that training SAE-V model for MLLMs with multimodal data is effective for interpreting MLLMs, and even LLMs, as SAE model trained solely on textual data fail to extract and disentangle the hidden representations of MLLMs effectively. Moreover, SAE-V model demonstrates superior capability in reconstructing the reasoning features of MLLMs compared to the standard SAE model.
3.2 Apply SAE-V Model on Multimodal Data
Figure 6: Case analysis of image patch filtering using $\mathcal{L}_{0}$ metric. We rank and filter image patches according to the number of features activated on them. The top row shows the original image (a) and its reduced-patch versions retaining 75% (b), 50% (c), and 25% (d). In this dog image, the patches are filterd out from edge to the middle and preserved almost only dog patches, suggesting that SAE-V model is preserving the main semantic information of the image.
In this section, we conduct an image classification task on the ImageNet dataset (Russakovsky et al., 2015) to investigate whether SAE-V can capture the key information within images and to validate the effectiveness of the methods proposed in Section 2.2 on multimodal data. We apply 4 methods, namely $\mathcal{L}_{0}$, $\mathcal{L}_{1}$, co-occurring $\mathcal{L}_{0}$, and cosine similarity score, where co-occurring $\mathcal{L}_{0}$ is defined as the number of features activated on at least one text and image token. The cosine similarity score is defined as the sum of cross-modal weights of features, consistent with Algorithm 1. We adopt these metrics to filter image patches, thus obtaining images that preserve 75%, 50%, and 25% patches, respectively.  [^3] We present complete algorithms of 4 methods in Appendix C.1.
Figure 7: The classification performance on ImageNet. We compare the classification accuracy after filtering the image patches with $\mathcal{L}_{0}$ method, $\mathcal{L}_{1}$ method, co-occurrence $\mathcal{L}_{0}$ method, cosine similarity score method, and the random baseline. All methods achieve high accuracy when preserving 75% or 50% patches and $\mathcal{L}_{0}$ method, $\mathcal{L}_{1}$ method, and cosine similarity score method maintains high accuracy even in the least patches. The result shows that SAE-V is efficient in capturing critical information from images.
Case Analysis
Figure 6 illustrates the image patches when using $\mathcal{L}_{0}$ metric for filtering. Even when employing the simplest $\mathcal{L}_{0}$ metric, SAE-V is still able to effectively capture the critical semantic information of the image.  [^4] More cases and analyses are presented in Appendix C.2.
Quantized Results
The quantized results are presented in Figure 7, where we observe that, for all methods, preserving 75% or 50% of the patches achieves an accuracy close to that obtained using the full image. In the challenging scenario where only 25% of the patches are retained, $\mathcal{L}_{0}$ method, $\mathcal{L}_{1}$ method, and cosine similarity score method maintain accuracy levels close to 70%, and all methods significantly surpass the accuracy obtained by the random preservation method. These results demonstrate that SAE-V can accurately capture critical information in images and that the methods proposed in Section 2.2 effectively utilize SAE-V features during inference.  [^5] More experiments and their quantative analysis are presented in Appendix  C.3
4 Alignment Experiment
In this section, we adopt cosine similarity score rank- ing algorithm as a data filter (as shown in Algorithm 1) to acquire high-quality data for model alignment.
4.1 Experiment Setup
(a) Performance on Align-Anything dataset
(b) Performance on MMInstruct dataset
(c) Performance on RLAIF-V dataset
Figure 8: Ablation study of SAE-V-based data filter method on Chameleon-7B model and different datasets. We replicate SAE-V-based data filter on Chameleon-7B model and three distinct datasets. The results show that SAE-V-based data filter method is still effective on different datasets and different model architectures and image encoding methods other than CLIP. (a) On Align-Anything dataset, both cosine similarity and co-occurrence filters perform better than random filters at almost every data percentage, and the cosine similarity filter achieved a score of 52.1 (120% of the full dataset’s performance) with 70% data. (b) On MMInstruct, the cosine similarity filter outperforms the random filter baseline at every data proportions. Specifically, the cosine similarity filter achieved a score of 61.2 (117% of the full dataset’s performance) with 60% data. (c) On RLAIF-V dataset, the cosine similarity filter achieves the highest score of 54.2 (125% of the full dataset’s performance) with only 10% data, demonstrating its supreme efficiency.
Dataset and Model
Consistent with Section 3.1.1, we selected LLaVA-NeXT-7B (Liu et al., 2024a) and Chameleon-7B (Team, 2024) for our alignment experiment. Since the LLaVA-NeXT-7B model is rather powerful in multimodal capabilities, we selected the Align-Anything (Ji et al., 2024) text-image-to-text dataset for our experiment. Align-Anything is a 400K multimodal preference dataset containing fine-grained annotated multimodal input-output preference data, and we used the 40K subset of text-image input and text output in our experiment.
Algorithm
We adopt the cosine similarity score ranking algorithm (shown in Algorithm 1) as a filter to exclude data with low scores. In addition, we also adopt two algorithms, the $\mathcal{L}_{0}$ ranking, and the co-occurrence ranking.  [^6] Detailed descriptions of these ablation algorithms and their corresponding hyperparameters are provided in Appendix B.
Evaluation
To evaluate the efficiency of our methods, we applied Direct Preference Optimization (DPO) to the model using the filtered datasets.  [^7] We present detailed training parameters in Appendix B.2. We then evaluate the multimodal capabilities of the model using LLaVA-Bench (Liu et al., 2024a) benchmarks.
Figure 9: The performance of SAE-V-based data filter method We evaluated SAE-V-based data filter method on LLaVA-NeXT-7B model, Align-Anything dataset, and LLaVA-Bench benchmark. The results show that all SAE-V-based methods significantly outperform the random selection baseline, while the cosine similarity filter achieved 108% of the full dataset’s performance with only 20% of the data, and the co-occurrence filter peaked at 50% of the data, reaching a score of 108.17. As a more straightforward utilization of SAE-V model, the L0 filter also generally outperforms the random selection baseline.
4.2 Experiment Results
We performed SAE-V-based data filter with different filtering ratios on the LLaVA-NeXT-7B model and Align-Anything dataset. The filtered datasets were then used to fine-tune MLLMs, which were evaluated on LLaVA-Bench. The results (shown in Figure 9) demonstrate that our SAE-V-based filtering method effectively enhances the alignment of LLaVA-NeXT-7B, even with reduced data. Since most of the data in Align-Anything contribute positively to model alignment, the performance of the model is higher than the base model without any fine-tuning in most cases. At any data filter proportion, the SAE-V-based data filtering method outperforms the random selection baseline, with the best result being 108.17 (115% of the full dataset’s performance) achieved using 50% filtered data from the cooccurrence filter, and 104.20 (108% of the full dataset’s performance) achieved using 20% filtered data from the cosine similarity filter. However, as the dataset inevitably contains some low-quality data, the performance is optimal with a moderate data proportion and shows a downward trend as the data proportion increases.
4.3 Relationship between MLLM Capability and SAE-V Features
Figure 10: The relationship between average cosine similarity score and MLLM performance. We measure the average cosine similarity score of models in Section 4.2, and fit a linear relationship between model performance and average cosine similarity score. The correlation coefficient of the correlation reaches 0.84, suggesting that higher similarity scores on SAE-V features correspond to enhanced MLLM performance.
In the previous section, we demonstrated the effectiveness of utilizing the cosine similarity score for data filters in model training. To further investigate the relationship between model performance and cross-modal similarity, as measured by cosine similarity of SAE-V features, we further measure the average cosine similarity score of these models. Given a dataset, we apply the cosine similarity score ranking algorithm (shown in Algorithm 1) to the MLLM, and we define the MLLM’s average cosine similarity score as the mean score of all non-zero cross-modal weight SAE-V features.
We calculated the average cosine similarity scores for the models discussed in Section 4.2. The result (shown in Figure 10) revealed a positive correlation between the average cosine similarity score of SAE-V feature and the performance of MLLM, suggesting that higher similarity scores of SAE-V features correspond to enhanced MLLM performance.
Method 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%
SAE-V 94.2 98.2 106.8 114.9 114.5 114.8 112.9 112.3 111.0 109.5 98.5
Random 94.2 96.5 97.0 98.3 95.3 93.7 96.5 98.8 98.2 96.8 98.5
Table 2: The performance of SAE-V-based data filter method on 1/20 Align-Anything dataset The performance of SAE-V-based data filter method on 1/20 Align-Anything dataset closely resembles that in Figure 9, confirming that alignment metrics on a small validation set resemble the distribution on the complete dataset, enabling efficient hyperparameter selection.
Method Performance
0% 20% 40% 60% 80% 100%
SAE-V 104.6 105.8 116.7 112.3 112.0 111.2
Random 104.6 105.3 105.8 107.2 110.4 111.2
Table 3: The performance of SAE-V-based data filter method on LLaVA-NeXT-Vicuna-13B We replicate SAE-V-based data filter on LLaVA-NeXT-Vicuna-13B and Align-Anything dataset. The result shows that SAE-V-based data filter outperforms the random selection baseline, and reached the highest performance of 116.67 with 40% data. This further indicates that SAE-V-based data filter could work regardless of model architecture and sizes.
4.4 Ablation Study
Ablation on Models
To prove that SAE-V-based data filter method could generalize to distinct model architectures, we replicate SAE-V-based data filter on Chameleon-7B and LLaVA-NeXT-Vicuna-13B models with the Align-Anything dataset. The result (shown in Figure 8 (a) and Table 3) demonstrates that SAE-V-based filter method still shows its effectiveness regardless of model architecture, sizes, and image encoding methods.
When using a smaller data proportion, the performance is strongly correlated with the data quantity, and thus, the differences between methods are minimal. However, with a larger data proportion, the SAE-V-based filter method significantly surpasses the random filter, achieving a peak of 52.13 (120% of the performance on full dataset) with 70% data on Chameleon-7B and 116.67 (104% of the performance on full dataset) with 40% data onLLaVA-NeXT-Vicuna-13B. The largest performance gap is observed in the 40-80% data range, while the differences converge again as the data proportion approaches 100%. This proves that SAE-V-based data filter is effective on architectures other than CLIP-based MLLM, and shows its potential to generalize across a wide range of models.
Ablation on Datasets
We also performed an ablation study on the datasets to be filtered. Since the LLaVA-NeXT-7B model is highly capable, most datasets fail to further enhance its multimodal abilities. Therefore, we selected the RLAIF-V (Yu et al., 2024) and MMInstruct (Liu et al., 2024b) datasets with the relatively weaker Chameleon-7B model for the dataset ablation study. The results (shown in Figure 8 (b) and Figure 8 (c)) further confirm that SAE-V-based data filter is working across different datasets. Moreover, on RLAIF-V, the cosine similarity filter could achieve a score of 54.23 (125% of the full dataset’s performance) by using only 10% of the data, and on MMInstruct, the cosine similarity filter could achieve a score of 61.2 (117% of the full dataset’s performance) with 60% data, demonstrating exceptional efficiency.
Ablation on Dataset Sizes
On real deployment scenarios, due to the large size of datasets, it is often difficult to use the SAE-V-based data filter method to search for filtering ratio parameters on the complete dataset. Therefore, we reproduced the SAE-V-based data filter method on a 1/20 subset of the Align-Anything dataset to demonstrate that in practical deployment situations, it is feasible to search for filter ratios on small-scale dataset subsets for application to the complete dataset. The results (shown in Table 2) indicate that the performance of SAE-V-based data filter on the small-scale dataset exhibits extremely similar trends with respect to filtered data ratios as those observed on the complete dataset (shown in Figure 9), which demonstrates that searching for filtered data ratios through small-scale dataset subsets is feasible.
Figure 11: The performance of SAE-V-based data filter and IFD metrics We replicate the IFD metric on MLLMs and compare the result with our cosine similarity filter. The results show that although the peak performance of SAE-V-based data filtering is a little lower compared to IFD metric, our method could achieve a generally similar performance compared to IFD metric without introducing additional models and training process.
Comparation with Other Filtering Methods
To validate the effectiveness of our SAE-V-based data filtering method, we conducted an ablation study comparing it with other similar data filtering approaches. Since there are currently no widely recognized data filtering methods specifically designed for multimodal data, we adapted the IFD metric (Li et al., 2024a) method to the multimodal setting. The result (shown in Figure 11) suggests that our data filter method achieves a performance comparable to the IFD metric. However, considering that the IFD metric needs to train an additional cherry model, our SAE-V-based data filter could directly fit various datasets, demonstrating greater generalizability and efficiency.
5 Related Work
Multimodal Large Language Model
MLLM is a type of LLM integrated with multimodal modules that incorporate multimodal information to deal with multimodal tasks. Based on the method of integrating vision features into the model, most MLLMs can be categorized into three types:
  • CLIP-based MLLMs: These models encode images with CLIP (Radford et al., 2021) and use MLP to project visual features. Examples include LLaVA (Liu et al., 2024a) series and NExT-GPT (Wu et al., 2024b).
  • Early-Fusion MLLMs: These models directly tokenize visual features for input. Examples include Chameleon (Team, 2024) and Janus (Wu et al., 2024a) series.
  • Q-Former-based MLLMs: These models use a structure similar to Q-Former (Li et al., 2023) to extract visual representations, represented by Qwen-VL (Bai et al., 2023) and MiniGPT-4 (Zhu et al., 2024).
Our study focuses on the CLIP-based and early-fusion MLLMs. Specifically, we select LLaVA-NeXT-7B and Chameleon-7B as the target models.
Mechanistic interpretability with Sparse Autoencoder
Mechanistic interpretability seeks to uncover and explain the internal mechanisms that enable models to understand input data and generate responses (Rai et al., 2024). Specifically, most current mechanistic interpretability methods focus on analyzing features, smaller units that contribute to performing explainable semantic tasks, within models (Olah et al., 2020).
Sparse Autoencoder (SAE) aims to learn sparse and interpretable features from polysemantic model representations (Yun et al., 2021; Bricken et al., 2023; Sharkey et al., 2022; Peigné, 2023; Elhage et al., 2022). By introducing sparsity constraints, the activation values in the hidden layers of SAE are mostly zero, allowing SAE to encode polysemantic features in LLM to monosemantic ones. Zhang et al. (2024) firstly attempted to apply SAE to analyze the open-semantic features of MLLMs.
In this paper, we extended the scope of SAE to MLLMs, thereby building SAE-V. We further demonstrated SAE-V’s capability and transferability on MLLMs, and built a data filter tool based on SAE-V to enhance multimodal alignment.
Data Filter in Alignment
Data filtering ensures that only relevant high-quality data are used during the alignment of LLMs or MLLMs, thus reducing the quantity of data while achieving greater performance (Zhou et al., 2023; Chen et al., 2024; Du et al., 2023; Li et al., 2024b, a; Tu et al., 2025). For example, LIMA (Zhou et al., 2023), ALPAGASUS (Chen et al., 2024), and IFD (Li et al., 2024a) use human annotation, API annotation and train a new model for annotation to score data separately. Our method, SAE-V-based data filter, provides a self-guided and interpretable metric to evaluate the similarity of multimodal data, which indicates their qualities. The method is stable and efficient for models of different architectures.
6 Conclusion
This work introduced SAE-V, a framework that extends SAE to MLLMs and improves their alignment. Through experiments on LLaVA-NeXT-7B and Chameleon-7B, we demonstrated that SAE-V model demonstrates excellent capability and transferability in interpreting MLLM and multimodal data, and SAE-V-based data filtering methods could achieve more than 110% performance with less than 50% data. These results highlight SAE-V’s potential to enhance multimodal model interpretability and alignment efficiently.
Limitation
While SAE-V introduces significant advancements in interpreting multimodal models and enhancing alignment through mechanistic analysis, several limitations remain unaddressed and warrant further exploration: (1) Although SAE-V demonstrates superior interpretability and data filtering efficiency compared to SAE, the theory behind SAE-V, especially the mathematical relationship between image-text similarity metrics, cross-modal co-occurrence features, and model performance, is not fully revealed. (2) Due to resource constraints, SAE-V is primarily evaluated on text and vision modalities, leaving its effectiveness on other modalities such as audio, video, and embodied AI systems unexplored. Our future work will focus on establishing a comprehensive theoretical foundation for SAE-V and extending its application to additional modalities, such as audio, video, and embodied AI systems, to broaden its utility and impact.
Impact Statement
The source code and checkpoints of SAE-V mentioned in this paper will be released under the CC BY-NC 4.0 license. This research has several potential risks that must be considered. The interpretability tools introduced in this work, while beneficial for alignment, could also be leveraged to manipulate or reverse-engineer model behaviors in unintended ways. Additionally, while SAE-V provides a self-guided filtering mechanism, it remains dependent on the initial dataset quality, meaning biases in the dataset could still propagate into the final model. We strongly condemn any malicious use of the SAE-V code and checkpoints and advocate for its responsible and ethical use.
Acknowledgement
This work is sponsored by National Natural Science Foundation of China (62376013, 623B2003).
Appendix A Details of Interpretability Experiment
A.1 Hyperparameter of SAE and SAE-V Models Training
Hyper-parameters SAE and SAE-V of LLaVA-NeXT/Mistral SAE and SAE-V of Chameleon/Anole
Training Parameters
total training steps 30000 30000
batch size 4096 4096
LR 5e-5 5e-5
LR warmup steps 1500 1500
LR decay steps 6000 6000
adam beta1 0.9 0.9
adam beta2 0.999 0.999
LR scheduler name constant constant
LR coefficient 5 5
seed 42 42
dtype float32 float32
buffer batches num 32 64
store batch size prompts 4 16
feature sampling window 1000 1000
dead feature window 1000 1000
dead feature threshold 1e-4 1e-4
SAE and SAE-V Parameters
hook layer 16 8
input dimension 4096 4096
expansion factor 16 32
feature number 65536 131072
context size 4096 2048
Table 4: Hyperparameters of training SAE and SAE-V models.
The hyperparameters of the training SAE and SAE-V are shown in Table 4. The differences in training parameters arise because the LLaVA-NeXT-7B model requires more GPU memory to handle vision input, so fewer batches can be cached. For the SAE and SAE-V parameters, we set different hook layers and context sizes based on the distinct architectures of the two models. We also experimented with different feature numbers on both models, but found that only around 30,000 features are actually activated during training. All training runs were conducted until convergence. All SAE and SAE-V training is performed on 8×A800 GPUs and each training typically takes around $21$ hours. We ensured that the variations in the parameters did not affect the experiment results.
Appendix B Details of alignment experiment
We present details of alignment experiment in this section, including algorithms and hyperparameters of algorithms and model training.
B.1 Algorithms in alignment experiment
The complete algorithm of $\mathcal{L}_{0}$-based Ranking and co-occurring $\mathcal{L}_{0}$-based Ranking are shown in Algorithm 2 and Algorithm 3. These two algorithms serve as ablation variants of the cosine similarity score Ranking (shown in Algorithm 1). The $\mathcal{L}_{0}$-based Ranking represents a straightforward algorithm that selects data by directly computing the sum of $\mathcal{L}_{0}$ for each data point. The co-occurring $\mathcal{L}_{0}$-based Ranking takes an initial step toward cross-modal consideration by only counting features that are activated across both modalities. Building upon these algorithms, we further developed the cosine similarity score Ranking approach.
Algorithm 2 $\mathcal{L}_{0}$-based Ranking
  Require: multimodal dataset $\mathcal{D}=\{\bm{d}_{i}\}_{i=1}^{p}$; MLLM $\mathcal{M_{\theta}}$; SAE-V $\mathcal{S_{\theta}}$; features of SAE-V; activation bound: $\delta$$\mathcal{F_{\theta}}\text{:}\ \{\bm{f}_{k}\}_{k=1}^{n}$;
  Ensure: Ranked data $\mathcal{D}_{R}$
  for each $\bm{d}_{i}\in\mathcal{D}$ do
     $Z_{i}\leftarrow\mathcal{S_{\theta}}(\mathcal{M_{\theta}}(\bm{d}_{i}))$
     $F_{i}\leftarrow\{\bm{f}_{k}\text{: }\exists\ z_{j}\in Z_{i},\ \mathrm{s.t.}\ z_{j}\ \text{activates }\bm{f}_{k}\}$
     $\mathcal{L}_{0,i}\leftarrow|F_{i}|$
  end for
  $\mathcal{D}_{R}\leftarrow\textit{Sort}(\mathcal{D},\{\mathcal{L}_{0,i}\}_{i=1}^{n})$
Algorithm 3 Co-ocurring $\mathcal{L}_{0}$-based Ranking
  Require: Text token vocabulary: $\mathcal{T}$; vision token vocabulary: $\mathcal{V}$; multimodal dataset $\mathcal{D}=\{\bm{d}_{i}\}_{i=1}^{p}$; MLLM $\mathcal{M_{\theta}}$; SAE-V $\mathcal{S_{\theta}}$; features of SAE-V $\mathcal{F_{\theta}}\text{:}\ \{\bm{f}_{k}\}_{k=1}^{n}$; activation bound: $\delta$;
  Ensure: Ranked data $\mathcal{D}_{R}$
  Initialize coocurrence feature set of data $F_{i}\leftarrow\varnothing$
  for each $\bm{d}_{i}\in\mathcal{D}$ do
     Initialize activated token set of features $\mathcal{A}_{k}\leftarrow\varnothing$
     $H_{i}\leftarrow\mathcal{M_{\theta}}(\bm{d}_{i})$
     $Z_{i}\leftarrow\mathcal{S_{\theta}}(H_{i})$
     for each $\bm{f}_{k}\in\mathcal{F}_{\theta}$ do
        $\mathcal{A}_{k}\leftarrow\mathcal{A}_{k}\cup\{\bm{h}_{j}\text{:}\ \bm{h}_{j}\in H_{i},\bm{z}_{j}=\bm{e}_{j}Z_{i},\bm{z}_{jk}>\delta\}$
        if $\mathcal{A}_{k}\cap\mathcal{T}\neq\varnothing\land\mathcal{A}_{k}\cap\mathcal{V}\neq\varnothing$ then
           $F_{i}\leftarrow F_{i}\cup\{\bm{f}_{k}\}$
        end if
     end for
  end for
  $\mathcal{D}_{R}\leftarrow\textit{Sort}(\mathcal{D},\{|F_{i}|\}_{i=1}^{n})$
Hyperparameters of Algorithms 1,2,3
The hyperparameters of Algorithms 1,2,3 are shown in Table 5. We ensure that all parameters are the same to ensure a fair comparison between the algorithms.
Hyper-parameters Cosine similarity Coocurrence $\mathcal{L}_{0}$
top-K 5 5 5
text token vocabulary size 32000 32000 32000
vision token vocabulary size 64 64 64
activation bound 1 1 1
sample data size 1000 1000 1000
Table 5: Hyper-parameters of Algorithm 1,2,3.
B.2 Hyperparameter of Model Training
In this section, we list out the hyperparameters used for model training through SFT and DPO (shown in Table 6). All SAE training is performed on 8×A800 GPUs. To ensure fair comparison between algorithms, we maintained consistent parameter settings across all experiments.
Hyper-parameters SFT DPO
max length 4096 4096
per device train batch size 8 8
per device eval batch size 8 8
gradient accumulation steps 4 4
LR scheduler type cosine cosine
LR 1e-6 1e-6
warmup steps 10 10
eval steps 50 50
epochs 3 3
val size 0.1 0.1
bf16 True True
Table 6: Hyperparameters of SFT training and DPO training.
Appendix C Details of Applying SAE-V on Multimodal Data
In this section, we present implementation details of the SAE-V application experiments. We enumerate 4 image patch selection algorithms employed in this study and provide additional case analyses. These comprehensive results further demonstrate the robust inference capabilities of SAE-V.
C.1 Algorithm
The complete algorithms of $\mathcal{L}_{0}$, $\mathcal{L}_{1}$, co-occurring $\mathcal{L}_{0}$, and cosine similarity score methods are shown in Algorithm 4, Algorithm 5, Algorithm 6 and Algorithm 7.
Algorithm 4 $\mathcal{L}_{0}$ patch filter
  Require: Vision token vocabulary: $\mathcal{V}$; image $V$; fixed Prompt $T$; MLLM $\mathcal{M_{\theta}}$; SAE-V $\mathcal{S_{\theta}}$; features of SAE-V $\mathcal{F_{\theta}}\text{:}\{\bm{f}_{k}\}_{k=1}^{n}$; activation bound $\delta$; mask rate $\gamma$;
  Ensure: Filtered image $V^{\prime}$
  Initialize score of each patch $p_{i}\leftarrow 0$
  $H\leftarrow\mathcal{M}_{\theta}(T,V)$
  $Z\leftarrow\mathcal{S}_{\theta}(H)$
  for each $\bm{h}_{i}\in H$ do
     if $\bm{h}_{i}\in\mathcal{V}$ then
        $p_{i}=\sum_{j}\mathbf{1}(z_{ij}>\delta)$
     end if
  end for
  $K\leftarrow\lfloor\gamma|I|\rfloor$
  $V^{\prime}\leftarrow\textit{TopK}(v_{i}\in V)\text{ sorted by }p_{i}$
Algorithm 5 $\mathcal{L}_{1}$ patch filter
  Require: Vision token vocabulary: $\mathcal{V}$; image $V$; fixed Prompt $T$; MLLM $\mathcal{M_{\theta}}$; SAE-V $\mathcal{S_{\theta}}$; features of SAE-V $\mathcal{F_{\theta}}\text{:}\{\bm{f}_{k}\}_{k=1}^{n}$; activation bound $\delta$; mask rate $\gamma$;
  Ensure: Filtered image $V^{\prime}$
  Initialize score of each patch $p_{i}\leftarrow 0$
  $H\leftarrow\mathcal{M}_{\theta}(T,V)$
  $Z\leftarrow\mathcal{S}_{\theta}(H)$
  for each $\bm{h}_{i}\in H$ do
     if $\bm{h}_{i}\in\mathcal{V}$ then
        $p_{i}=\sum_{j}(z_{ij})$
     end if
  end for
  $K\leftarrow\lfloor\gamma|I|\rfloor$
  $V^{\prime}\leftarrow\textit{TopK}(v_{i}\in V)\text{ sorted by }p_{i}$
These algorithms take images as input and produce masked images, where the masking proportion is determined by the mask rate $\gamma$. All algorithms utilize the activation patterns of SAE-V features for patch filtering, with their primary distinctions lying in their methods of computing feature activation ($\mathcal{L}_{0}$, $\mathcal{L}_{1}$) and measuring cross-modal similarity (co-occurring $\mathcal{L}_{0}$, cosine similarity score).
[…]
  Require: Text token vocabulary $\mathcal{T}$; vision token vocabulary: $\mathcal{V}$; image $V$; fixed Prompt $T$; MLLM $\mathcal{M_{\theta}}$; SAE-V $\mathcal{S_{\theta}}$; features of SAE-V $\mathcal{F_{\theta}}\text{:}\{\bm{f}_{j}\}_{j=1}^{n}$; activation bound $\delta$; mask rate $\gamma$;
  Ensure: Filtered image $V^{\prime}$
  Initialize score of each patch $p_{i}\leftarrow 0$, co-occuring feature set $F\leftarrow\varnothing\ \text{and activated token set of features }\mathcal{A}_{j}\leftarrow\varnothing$
  $H\leftarrow\mathcal{M}_{\theta}(T,V)$
  $Z\leftarrow\mathcal{S}_{\theta}(H)$
  for each $\bm{f}_{j}\in\mathcal{F}_{\theta}$ do
     $\mathcal{A}_{j}\leftarrow\mathcal{A}_{j}\cup\{\bm{h}_{i}\text{:}\ \bm{h}_{i}\in H,\bm{z}_{i}=\bm{e}_{i}Z,\bm{z}_{ij}>\delta\}$
     if $\mathcal{A}_{j}\cap\mathcal{T}\neq\varnothing\land\mathcal{A}_{j}\cap\mathcal{V}\neq\varnothing$ then
        $F\leftarrow F\cup\{\bm{f}_{j}\}$
     end if
  end for
  for each $\bm{h}_{i}\in H$ do
     if $\bm{h}_{i}\in\mathcal{V}$ then
        $p_{i}=\sum_{j}\mathbf{1}(z_{ij}>\delta\ \land\ \bm{f}_{j}\in F)$
     end if
  end for
  $K\leftarrow\lfloor\gamma|I|\rfloor$
  $V^{\prime}\leftarrow\textit{TopK}(v_{i}\in V)\text{ sorted by }p_{i}$
Algorithm 7 Cosine similarity score patch filter
  Require: Text token vocabulary $\mathcal{T}$; vision token vocabulary: $\mathcal{V}$; image $V$; fixed Prompt $T$; MLLM $\mathcal{M_{\theta}}$; SAE-V $\mathcal{S_{\theta}}$; features of SAE-V $\mathcal{F_{\theta}}\text{:}\{\bm{f}_{j}\}_{j=1}^{n}$; activation bound $\delta$; mask rate $\gamma$;cosine similarity weight $\{\omega_{j}\}_{j=1}^{n}$
  Ensure: Filtered image $V^{\prime}$;
  Initialize score of each patch $p_{i}\leftarrow 0$, co-occuring feature set $F\leftarrow\varnothing\ \text{and activated token set of features }\mathcal{A}_{j}\leftarrow\varnothing$
  $H\leftarrow\mathcal{M}_{\theta}(T,V)$
  $Z\leftarrow\mathcal{S}_{\theta}(H)$
  for each $\bm{f}_{j}\in\mathcal{F}_{\theta}$ do
     $\mathcal{A}_{j}\leftarrow\mathcal{A}_{j}\cup\{\bm{h}_{i}\text{:}\ \bm{h}_{i}\in H,\bm{z}_{i}=\bm{e}_{i}Z,\bm{z}_{ij}>\delta\}$
     if $\mathcal{A}_{j}\cap\mathcal{T}\neq\varnothing\land\mathcal{A}_{j}\cap\mathcal{V}\neq\varnothing$ then
        $F\leftarrow F\cup\{\bm{f}_{j}\}$
     end if
  end for
  for each $\bm{h}_{i}\in H$ do
     if $\bm{h}_{i}\in\mathcal{V}$ then
        $p_{i}=\sum_{j}\mathbf{1}(z_{ij}>\delta\ \land\ \bm{f}_{j}\in F)\ \omega_{j}$
     end if
  end for
  $K\leftarrow\lfloor\gamma|I|\rfloor$
  $V^{\prime}\leftarrow\textit{TopK}(v_{i}\in V)\text{ sorted by }p_{i}$
C.2 Case Analysis
We present 4 cases in Figure 12, corresponding to each of our metric in Section 3.2. The cases intuitively show that $\mathcal{L}_{0}$ method and cosine similarity score method are more capable of identifying significant patches in images compared to other methods, which aligns with the quantized results shown in Figure 7.
Figure 12: Case Analysis of all image patch filtering methods in Section 3.2. We present the original image (a) and 4 case for methods, $\mathcal{L}_{0}$ (b), $\mathcal{L}_{1}$ (c), co-occurring $\mathcal{L}_{0}$ (d) and cosine similarity score (e). Each case contains 3 images as preserving top 75% patches, top 50% patches and top 25% patches.
We present 5 cases filtered with the cosine similarity score method in Figure 13. The results show that SAE-V model performs excellently in capturing critical patches in images.
Figure 13: Case Analysis of cosine similarity score method in Section 3.2. 5 cases filtered with cosine similarity score method are shown in the Figure. Each case contains contains 4 images as original image, preserving top 75% patches, top 50% patches and top 25% patches.
C.3 Quantative Analysis
Masking (%) 0 25 50 75
Text Accuracy (%) 80.2 77.7 66.5 53.3
Image Accuracy (%) 80.2 78.8 78.4 70.7
Table 7: Applying SAE-V on VQA tasks We apply SAE-V-based patch filtering experiment on the text and image of VQA task separately, using A-OKVQA validation set with LLaVA-NeXT-7B model. The experiment shows that image information demonstrates a lower compression rate (more redundancy) than text. Moreover, text masking shows a roughly linear relation of accuracy and masked token, while image filter maintains performance until 50%, with a more significant drop only appearing at 75%, suggesting that the information of text is more evenly distributed compared to image.
We performed additional experiments on SAE-V-based patch filtering. To be specific, we use SAE-V-based patch filter on the text and image part of each VQA question separately, and test the accuracy when part of the information is masked. The results show that image information demonstrates a lower compression rate (more redundancy) than text. Moreover, text masking shows a roughly linear relation of accuracy and masked token, while image filter maintains performance until 50%, with a more significant drop only appearing at 75%, suggesting that the information of text is more evenly distributed compared to image.